A single circular loop of radius 0.23m carries a current of 2.7A in a magnetic field of 0.96T .

Part A
What is the maximum torque exerted on this loop?
Answer #1

torque max = pi*0.23^2*2.7*0.96 = IAB

= 0.43 Nm
Answer #2

For a single loop of area A, with current I, in a field B, the maximum torque is given by

T = A x I x B

But first we need to calculate A.

A = pi x R^2  

T = pi x R^2 x I x B


T = 3.14 x (0.23)^2 x (2.7) x (0.96) = 0.4305 N-m (ANSWER)